Reports 1-1 of 1 Clear search Modify search
Injection system (General activities)
derossi, chiummo, gosselin - 18:09 Tuesday 21 September 2021 (53212) Print this report
Laser power injected into the ITF increased to 40 W

This morning we increased the pumping current of the fiber amplifier in order to reach 40 W injected in the interferometer (plot 1).

We then blocked the beam to allow the work in the mode cleaner building. We relocked the injection at around 16h30 LT and we measured the TF of the IMC.

We found that the attenuation on the locking electronics was at the minimum (1 dB) and the UGF at around 99 kHz. We tried to increase the attenuation but some oscillations appeared, we thus left the attenuation at 1 dB (plot 2). We realized that once relocked the injected power to the ITF had lowered to 39 W (plot 3 and 4).

We will let as it is for tonight and we will see tomorrow its evolution during the night.

Images attached to this report
Search Help